# Strategic Classification with Non-Linear Classifiers

### v1

@@ B.10 Proof of Theorem 3
$=O\big{(}k\,d^{2}\,log(k)\big{)}+O\big{(}k\,\,d\,\,log(k)\big{)}$
$=O\big{(}k\,d^{2}\,log(k)\big{)}$
B.11 Proof of Proposition 5
Proposition 5:
For all $d$, there exist (non-degenerate) distributions on $\mathbb{R}^{d}$ that are realizable in the standard setting, but whose strategic accuracy is the majority class rate, $\max_{y}p(y)$.
Proof.
Fix $d$ and choose any arbitrary set of positive points ${\cal{X}}_{1}$. When $d=1$, the set of negative points ${\cal{X}}_{0}$ can be constructed by placing a negative point $\delta$ to the left and right of each positive point. When $d>1$, this generalizes to selecting $d+1$ points around each positive point $x\in{\cal{X}}_{1}$ that form an $\mathbb{R}^{d}$ tetrahedron $T^{d}_{\delta}(x)$ [^7] We can assume without loss of generality that $\bigcap_{x_{1},x_{2}\in{\cal{X}}_{1}}T^{d}_{\delta}(x)=\varnothing\,\,\forall x_{1},x_{2}\in{\cal{X}}_{1}$ since there are infinite options for each $T^{d}_{\delta}$ and we can just pick ones that do not overlap.. In this case, the classifier $h(x)=\mathds{1}{\{{x\in{\cal{X}}_{1}}\}}$ achieves perfect standard accuracy, and the classifier that classifies all points as negative will achieve a strategic accuracy of $\frac{d+1}{d+2}$. Note that for each point $z$ in space to which $x\in{\cal{X}}_{1}$ can strategically move, there exists at least one $x^{\prime}\in T^{d}_{\delta}(x)$ that can as well. Therefore, for each positive point $x\in{\cal{X}}_{1}$ that achieves a correct positive classification by strategically moving to $z$ ($z$ can be the same as $x$ if it doesn’t move), there is a unique negative point $x^{\prime}\in T^{d}_{\delta}(x)$ that achieves an incorrect negative classification by strategically moving to $z$, so the strategic accuracy does not benefit from correctly classifying any positive points.
@@ B.12 Proof of Proposition 6
There exists a distribution which is realizable in the standard setting, but whose accuracy under any $h_{\Delta}$ is at most $0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$.
@@
Consider set of lattice points in $\mathbb{R}^{1}$ with alternating labels. Although $f(x)=sin(\pi x)$ achieves perfect standard accuracy on this distribution, we will show that the maximum strategic accuracy for any model class is $0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$, which approaches 0.5 as $\alpha\to\infty$. Note that in $R^{1}$, the minimum size of a positive interval is $2\alpha$. Additionally, for each positive interval $I$ of length $k$, $I$ will contain either $\lfloor k\rfloor$ or $\lfloor k\rfloor+1$ points and include at most one more positive point than negative point. Therefore, each interval $I$ will at best be correct on $\frac{\lfloor k\rfloor+1}{2}$ out of $\lfloor k\rfloor$ points, which is optimized for $k=2\alpha$ (i.e. the minimum possible value of $k$). Furthermore, because the negative intervals may also only have at most one more negative point than positive point, the negative intervals are optimized when they include only a single negative point. As such, an optimal strategic classifier $h$ consists of repeated positive intervals that includes $\lfloor 2\alpha\rfloor$ points followed by a negative interval that includes one negative point. For each positive and negative interval group, $h$ correctly classifies at most $\frac{\lfloor 2\alpha\rfloor+1}{2}+1$ out of $\lfloor 2\alpha+1\rfloor$ points correctly for an overall maximum accuracy of $\frac{\frac{\lfloor 2\alpha+1\rfloor}{2}+1}{\lfloor 2\alpha+1\rfloor}=0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$. ∎
@@ C.1 Partial Expansion
C.2 Merging/Splitting Regions
As seen in Fig. 9 (Right), two positive regions can merge together when they are separated by a narrow pass in $h$. If the narrow pass is part of a larger negative region, the region will be split in two, which can potentially increase the complexity of the decision boundary (for example in Fig. 9 (Left)).
C.3 Types of Impossible Effective Classifiers
As stated in Sec. 5.3, we identify four "types" of classifiers $g$ that cannot exists as effective classifiers. Formally, these classifiers are ones which include either:
@@ C.3 Types of Impossible Effective Classifiers
    A small continuous region of positively classified points $C$ where $\exists z\in R^{d}$ such that $x\in B_{\alpha}(z)\,\forall x\in C$.
@@
Let $x\in C$ be some point in a small positive region in $g$. For $x$ to be classified as positive, there must be a point $x^{\prime}\in B_{\alpha}(x)$ which is classified positively by $h$ and not reachable by any $z\in{\cal{X}}_{0}(h)$ (see requirements 1 and 2 in Appendix B.4). However, any point $x^{\prime}\in B_{\alpha}(x)\,\cap C$ is reachable by ${\cal{X}}_{0}(h)$ and any point $x^{\prime}\in B_{\alpha}(x)\,\cap C^{c}$ is also reachable by ${\cal{X}}_{0}(h)$ since it is closer to some $z\in{\cal{X}}_{0}(h)$ than to $x$. Therefore, there are no points $x^{\prime}\in B_{\alpha}(x)$ which are not reachable by any $z\in{\cal{X}}_{0}(h)$, so $g$ cannot be an effective classifier.
@@
We will show that any $g$ with directional curvature greater than $1/\alpha$ anywhere on the decision boundary will fail Prop. 3. In other words, if the directional curvature at point $x$ is greater than $1/\alpha$ , then the point $x+\alpha\hat{n}_{x}$ will be reachable by some point that is negatively classified by $g$. In order for the point $x+\alpha\hat{n}_{x}$ to be reachable by $x$ but not any point negatively classified by $g$, the ball $B_{\alpha}(x+\alpha\hat{n}_{x})$ – which is the set of points that can strategically move to $x+\alpha\hat{n}_{x}$ – must not intersect $g$ and must also be strictly tangent to it an $x$. However, if the curvature at point $x$ is greater than $1/\alpha$, then the maximum radius $r_{insc}$ of a ball that is strictly tangent to $x$ is less than $\alpha$ so $g$ does not meet Prop. 3.
In $\mathbb{R}^{2}$, Lemma 1 proves that the radius $r_{insc}$ of the maximum inscribed circle at $x$ is no bigger than the radius $r_{osc}=\frac{1}{\kappa}$ of the osculating circle at $x$. We therefore have that $r_{insc}\leq r_{osc}=\frac{1}{\kappa}<\alpha$. In higher dimensions (see Appendix A.1), there is no longer a single normal curvature. Instead, we define the set $K$ of curvatures in each plane $p\in P$ defined by the normal vector and some tangent vector at $x$. Note that $B_{\alpha}(x+\alpha\hat{n}_{x})$ is only strictly tangent to the decision boundary at $x$ if it’s intersection with each $p\in P$ (a circle) is strictly tangent at $x$ to the curve obtained by intersecting the decision boundary with the plane $p$. Since each of these cases are in $\mathbb{R}^{2}$, each one will only hold if the curvature at $x$ in $p$ is no more than than $1/\alpha$. Therefore, if $\kappa^{*}=\sup(K)>1/\alpha$ for any $x$ on the decision boundary, then the maximum radius of a ball that is strictly tangent to $x$ is less than $\alpha$, so $g$ does not meet Prop. 3.
@@
The set of points which are reachable by point $x$ but not point $x+\delta\vec{v}$ for unit vector $\vec{v}$ and $\delta\to 0$ all satisfy $\vec{v}\cdot x\leq 0$.
@@
Define an orthonormal basis $v_{1},v_{2}...v_{d}$ with $v_{1}=v$. For any point $y$ such that $\vec{v}\cdot y>0$, the vector connecting $x$ and $y$ can be written as $a_{1}v_{1}+a_{2}v_{2}...+a_{d}v_{d}$ while the vector connecting $x+\delta\vec{v}$ and $y$ can be written as $(a_{1}-\delta)v_{1}+a_{2}v_{2}...+a_{d}v_{d}$. Since $\delta<a_{1}$, $y$ is closer to $x+\delta\vec{v}$ than $x$. Therefore, any point that is reachable by $x$ and not by $x+\delta\vec{v}$ must satisfy $\vec{v}\cdot x\leq 0$. ∎
@@
As depicted in Fig. 10, let $L_{1}$ and $L_{2}$ be two $R^{d}$ hyperplane segments that intersect on the decision boundary of $g$ at the $R^{d-1}$ hyperplane segment $I$, and let $x$ be some point on I. Let $0<\theta<\pi$ be the angle between $L_{1}$ and $L_{2}$ on the positive side. Let $\vec{v}_{1}$ be the unit vector in $L_{1}$ that runs through $x$ and is orthogonal to $I$; $\vec{v}_{2}$ be the unit vector in $L_{2}$ that runs through $x$ and is orthogonal to $I$; and $\vec{v}_{C}=-\frac{\vec{v}_{1}+\vec{v}_{2}}{||\vec{v}_{1}+\vec{v}_{2}||}$. Note that for $\delta\to 0$ the points $x_{1}=x+\delta\vec{v}_{1}$ and $x_{2}=x+\delta\vec{v}_{2}$ are on the decision boundary while the point $x_{12}=x+\delta\vec{v}_{C}$ is classified as positive and $x_{C}=x-\delta\vec{v}_{C}$ is classified as negative. Additionally, if we define $\vec{v}_{A}$ as the the result of $\vec{v}_{1}$ being rotated $\epsilon<\frac{\pi-\theta}{2}$ around $I$ away from the positive region and $\vec{v}_{B}$ as the the result of $\vec{v}_{2}$ being rotated $\epsilon$ around $I$ away from the positive region, then both $x_{A}=x+\delta\vec{v}_{A}$ and $x_{B}=x+\delta\vec{v}_{B}$ are classified as negative.
@@
Assume that there exists a classifier $h$ associated with the effective classifier $g$. Therefore, there exists a point $y$ that is reachable to $x$ but not to $x_{A}$, $x_{B}$, or $x_{C}$ (which are classified as negative by $g$). By Lemma 4, any point $y$ that is reachable under strategic movement by point $x$ but not $x_{A}$, $x_{B}$, or $x_{C}$ must at least satisfy $y\cdot\vec{v}_{A}\leq 0,\text{ }y\cdot\vec{v}_{B}\leq 0,\text{ and }y\cdot\vec{v}_{C}\leq 0$. However, because $\vec{v}_{A}+\vec{v}_{B}$ points in the same direction as $\vec{v}_{1}+\vec{v}_{2}$ – and therefore the opposite direction of $\vec{v}_{C}$ – we must have that $y\cdot\vec{v}_{C}=0$ and $y\cdot(\vec{v}_{A}+\vec{v}_{B})=0$, which implies that both $y\cdot\vec{v}_{A}=0$ and $y\cdot\vec{v}_{B}=0$. However, this implies that the vector $\vec{y}$ is orthogonal to both $\vec{v}_{A}$ and $\vec{v}_{B}$, which can only happen if $y$ lies on $I$ or if the angle between $\vec{v}_{A}$ and $\vec{v}_{B}$ is either 00 or $\pi$. In the former case, $y$ is then reachable by the negatively classified point that is just over the decision boundary from it. The latter case is also not an option because the angle between $\vec{v}_{A}$ and $\vec{v}_{B}$ is $\theta_{AB}=\theta+2\epsilon$ and we have ensured that $0<\theta+2\epsilon<\theta+2\frac{\pi-\theta}{2}=\pi$. In either case, it is impossible for $g$ to classify $x$ as positive while classifying $x_{A}$, $x_{B}$, and $x_{C}$ as negative, so there is no $h$ such that $g=h_{\Delta}$. ∎
Figure 10: Diagram for proof of case 4.
C.4 Non-bijectiveness of Classifier Mapping

### v2

@@ B.11 Proof of Theorem 4
$=O\big(d^{2}\,k\,log(k)\big)+O\big(d\,\,k\,\,log(k)\big)$
$=O\big(d^{2}\,k\,log(k)\big)$
@@ B.12 Proof of Proposition 5
For all $d$, there exist (non-degenerate) distributions on $\mathbb{R}^{d}$ that are realizable in the standard setting, but whose maximum strategic accuracy is the majority class rate, $\max_{y}p(y)$.
@@
Fix $d$ and choose any arbitrary set of positive points ${\cal{X}}_{1}$. When $d=1$, the set of negative points ${\cal{X}}_{0}$ can be constructed by placing a negative point $\delta$ to the left and right of each positive point. When $d>1$, this generalizes to selecting $d+1$ points around each positive point $x\in{\cal{X}}_{1}$ that form an $\mathbb{R}^{d}$ simplex $T^{d}_{\delta}(x)$ [^7] We can assume without loss of generality that $\bigcap_{x_{1},x_{2}\in{\cal{X}}_{1}}T^{d}_{\delta}(x)=\varnothing\,\,\forall x_{1},x_{2}\in{\cal{X}}_{1}$ since there are infinite options for each $T^{d}_{\delta}$, and we can just pick ones that do not overlap.. In this case, the classifier $h(x)=\mathds{1}{\{{x\in{\cal{X}}_{1}}\}}$ achieves perfect standard accuracy, and the classifier that classifies all points as negative will achieve a strategic accuracy of $\frac{d+1}{d+2}$. Note that for each point q{q} in space to which $x\in{\cal{X}}_{1}$ can strategically move, there exists at least one $x^{\prime}\in T^{d}_{\delta}(x)$ that can as well. Therefore, for each positive point $x\in{\cal{X}}_{1}$ that achieves a correct positive classification by strategically moving to q{q} (q{q} can be the same as $x$ if it doesn’t move), there is a unique negative point $x^{\prime}\in T^{d}_{\delta}(x)$ that achieves an incorrect negative classification by strategically moving to q{q}, so the strategic accuracy does not benefit from correctly classifying any positive points. As a result, the maximum strategic accuracy is achieved by correctly classifying each negative point and incorrectly classifying each positive point, which gives the majority class rate.
B.13 Proof of Proposition 6
Proposition 6:
There exists a distribution that is realizable in the standard setting, but whose accuracy under any $h_{\Delta}$ is at most $0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$.
Proof.
Consider the set of lattice points in $\mathbb{R}^{1}$ with alternating labels. Although $f(x)=sin(\pi x)$ achieves perfect standard accuracy on this distribution, we will show that the maximum strategic accuracy for any model class is $0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$, which approaches 0.5 as $\alpha\to\infty$. Note that in $R^{1}$, the minimum length of a positive interval is $2\alpha$. Additionally, for each positive interval $I$ of length $k$, II will contain either $\lfloor k\rfloor$ or $\lfloor k\rfloor+1$ points and include at most one more positive point than negative points. Therefore, each interval $I$ will, at best, be correct on $\frac{\lfloor k\rfloor+1}{2}$ out of $\lfloor k\rfloor$ points, which is optimized for $k=2\alpha$ (i.e., the minimum possible value of $k$). Furthermore, because the negative intervals may also only have at most one more negative point than positive points, the negative intervals are optimized when they include only a single negative point. As such, an optimal strategic classifier $h$ consists of repeated positive intervals that include $\lfloor 2\alpha\rfloor$ points, followed by a negative interval that includes one negative point. For each positive and negative interval group, $h$ correctly classifies at most $\frac{\lfloor 2\alpha\rfloor+1}{2}+1$ out of $\lfloor 2\alpha+1\rfloor$ points correctly for an overall maximum accuracy of $\frac{\frac{\lfloor 2\alpha+1\rfloor}{2}+1}{\lfloor 2\alpha+1\rfloor}=0.5+\frac{1}{\lfloor 2\alpha+1\rfloor}$. ∎
@@ Appendix C Additional results
C.1 Other Cost functions
Our low-level analysis (Sec. 4) begins with a focus on the $\ell_{2}$-norm cost, both because it is the most popular choice in the literature, and because we found it to offer the best intuition. Nonetheless, most of our results hold more broadly, and in particular, all theorems and propositions in Sec. 5 generalize to any $\ell_{p}$ ($p\geq 1$) norm:
  • Prop. 4 applies generally, since negative-inside polytopes are still mapped to smaller negative-inside polytopes under general $\ell_{p}$ cost functions.
  • Thm. 1, Cor. 1, and the example in Fig. 4 of increasing VC dimension remain under minor tweaks to the constructions that take into account the differences in shapes of $\ell_{p}$ balls for different values of $p\geq 1$.
  • Thm. 2 holds for any cost function under which the maximum Euclidean distance any user can strategically move is bounded by a constant. This holds not only for $\ell_{p}$ norms, but also for most reasonable cost functions.
  • Thm. 3 is already stated for general $\ell_{p}$ norms.
  • Thm. 4 is stated for $\ell_{1}$, $\ell_{2}$, and $\ell_{\infty}$. However, a more general result can be stated with dependence on the VC dimension of the cost function balls (i.e., the set $C=\{c(x,x^{\prime})\leq\alpha\,\,\forall x\in\mathbb{R}^{d}\}$). The bound would then be $O(d^{2}\,k\,log(k)+\mathrm{VC}(C)\,)$, and would support any plug-in results for $\mathrm{VC}(C)$.
  • Prop. 5 and Prop. 6 also remain under minor tweaks to the constructions that take into account the differences in shapes of $\ell_{p}$ balls for different values of $p\geq 1$.
  • Obs. 3, 4, 5, and Cor. 2 hold for $\ell_{p}$ norms in general as well, since they are not shape dependent, but rather results of movement within a ball.
Thus, our claims regarding the fact that non-learnable classes can become learnable, the loss of universality, and the limits to approximation all hold more generally.
In terms of asymmetric norms such as the Mahalanobis norm, while these norms would complicate the curvature analysis (and subsequent impossibility results that build off them), all results in Sec. 5 would still hold. This is because while the shape of balls changes (like with other $\ell_{p}$ balls), fundamental properties like lines mapping to other lines do not. As such, the results in Sec. 5 would still hold for the exact same reasons as delineated above for $\ell_{p}$ balls.
As for feature-dependent (also known as “instance-wise”) costs, these are qualitatively different from the conventional global costs, in a way that can make the claims become irrelevant or even degenerate. For example, Sundaram et al. [33] show that on the one hand, even for linear $H$, instance-wise cost function can cause $\mathrm{VC}(H_{\Delta})=\infty$, and on the other hand, for separable costs (which are instance-wise) it holds that $\mathrm{VC}(H_{\Delta})\leq 2$ for any $H$.
@@ C.2 Partial Expansion
C.3 Merging/Splitting Regions
As seen in Fig. 9 (Right), two positive regions can merge together when they are separated by a narrow pass in $h$. If the narrow pass is part of a larger negative region, the region will be split into two, which can potentially increase the complexity of the decision boundary (for example, in Fig. 9 (Left)).
C.4 Types of Impossible Effective Classifiers
As stated in Sec. 4.4, we identify four "types" of classifiers $g$ that cannot exist as effective classifiers. Formally, these classifiers are ones which include either:
@@ C.4 Types of Impossible Effective Classifiers
    A small continuous region of positively classified points $C$ where $\exists{q}\in R^{d}$ such that $x\in B_{\alpha}({q})\,\forall x\in C$.
@@
Let ${x}\in C$ be some point in a small positive region in $g$. For x{x} to be classified as positive, there must be a point $x^{\prime}\in B_{\alpha}({x})$ which is classified positively by $h$ and not reachable by any ${q}\in{\cal{X}}_{0}(h)$ (see requirements 1 and 2 in Appendix B.4). However, any point $x^{\prime}\in B_{\alpha}({x})\,\cap C$ is reachable by ${\cal{X}}_{0}(h)$ and any point $x^{\prime}\in B_{\alpha}({x})\,\cap C^{c}$ is also reachable by ${\cal{X}}_{0}(h)$ since it is closer to some ${q}\in{\cal{X}}_{0}(h)$ than to x{x}. Therefore, there are no points $x^{\prime}\in B_{\alpha}({x})$ which are not reachable by any ${q}\in{\cal{X}}_{0}(h)$, so $g$ cannot be an effective classifier.
@@
We will show that any $g$ with directional curvature greater than $1/\alpha$ anywhere on the decision boundary will fail Prop. 3. In other words, if the directional curvature at point x{x} is greater than $1/\alpha$ , then the point ${x}+\alpha\hat{n}_{x}$ will be reachable by some point that is negatively classified by $g$. In order for the point ${x}+\alpha\hat{n}_{x}$ to be reachable by x{x} but not any point negatively classified by $g$, the ball $B_{\alpha}({x}+\alpha\hat{n}_{x})$ — which is the set of points that can strategically move to ${x}+\alpha\hat{n}_{x}$ — must not intersect $g$ and must also be strictly tangent to it at x{x}. However, if the curvature at point x{x} is greater than $1/\alpha$, then the maximum radius $r_{insc}$ of a ball that is strictly tangent to x{x} is less than $\alpha$, so $g$ does not meet Prop. 3.
In $\mathbb{R}^{2}$, Lemma 1 proves that the radius $r_{insc}$ of the maximum inscribed circle at x{x} is no bigger than the radius $r_{osc}=\frac{1}{\kappa}$ of the osculating circle at x{x}. We therefore have that $r_{insc}\leq r_{osc}=\frac{1}{\kappa}<\alpha$. In higher dimensions (see Appendix A.1), there is no longer a single normal curvature. Instead, we define the set $K$ of curvatures in each plane $p\in P$ defined by the normal vector and some tangent vector at x{x}. Note that $B_{\alpha}({x}+\alpha\hat{n}_{x})$ is only strictly tangent to the decision boundary at x{x} if it’s intersection with each $p\in P$ (a circle) is strictly tangent at x{x} to the curve obtained by intersecting the decision boundary with the plane $p$. Since each of these cases are in $\mathbb{R}^{2}$, each one will only hold if the curvature at x{x} in $p$ is no more than $1/\alpha$. Therefore, if $\kappa^{*}=\sup(K)>1/\alpha$ for any x{x} on the decision boundary, then the maximum radius of a ball that is strictly tangent to x{x} is less than $\alpha$, so $g$ does not meet Prop. 3.
@@
The set of points that are reachable by point $x$ but not point $x+\delta\vec{v}$ for unit vector $\vec{v}$ and $\delta\to 0$ all satisfy $\vec{v}\cdot x\leq 0$.
@@
Define an orthonormal basis $v_{1},v_{2}...v_{d}$ with $v_{1}=v$. For any point q{q} such that $\vec{v}\cdot{q}>0$, the vector connecting $x$ and q{q} can be written as $a_{1}v_{1}+a_{2}v_{2}...+a_{d}v_{d}$ while the vector connecting $x+\delta\vec{v}$ and q{q} can be written as $(a_{1}-\delta)v_{1}+a_{2}v_{2}...+a_{d}v_{d}$. Since $0<\delta<a_{1}$ (note that $\vec{v}\cdot{q}>0$ so $a_{1}>0$), q{q} is closer to $x+\delta\vec{v}$ than $x$. Therefore, any point that is reachable by $x$ and not by $x+\delta\vec{v}$ must satisfy $\vec{v}\cdot x\leq 0$. ∎
@@
[…]
@@
Assume that there exists a classifier $h$ associated with the effective classifier $g$. Therefore, there exists a point q{q} that is reachable by x{x} but not by $x_{A}$, $x_{B}$, or $x_{C}$ (which are classified as negative by $g$). By Lemma 4, any point q{q} that is reachable under strategic movement by point x{x} but not $x_{A}$, $x_{B}$, or $x_{C}$ must at least satisfy ${q}\cdot\vec{v}_{A}\leq 0,\text{ }{q}\cdot\vec{v}_{B}\leq 0,\text{ and }{q}\cdot\vec{v}_{C}\leq 0$. However, because $\vec{v}_{A}+\vec{v}_{B}$ points in the same direction as $\vec{v}_{1}+\vec{v}_{2}$ — and therefore the opposite direction of $\vec{v}_{C}$ — we must have that ${q}\cdot\vec{v}_{C}=0$ and ${q}\cdot(\vec{v}_{A}+\vec{v}_{B})=0$, which implies that both ${q}\cdot\vec{v}_{A}=0$ and ${q}\cdot\vec{v}_{B}=0$. However, this implies that the vector $\vec{{q}}$ is orthogonal to both $\vec{v}_{A}$ and $\vec{v}_{B}$, which can only happen if q{q} lies on $I$ or if the angle between $\vec{v}_{A}$ and $\vec{v}_{B}$ is either 0 or $\pi$. In the former case, q{q} is then reachable by the negatively classified point that is just over the decision boundary from it. The latter case is also not an option because the angle between $\vec{v}_{A}$ and $\vec{v}_{B}$ is $\theta_{AB}=\theta+2\epsilon$, and we have ensured that $0<\theta+2\epsilon<\theta+2\frac{\pi-\theta}{2}=\pi$. In either case, it is impossible for $g$ to classify x{x} as positive while classifying $x_{A}$, $x_{B}$, and $x_{C}$ as negative, so there is no $h$ such that $g=h_{\Delta}$. ∎
C.5 Non-bijectiveness of Classifier Mapping